Solve. Given a population of the scores of BSHS honor students in Mathematics as 1, 3, 4, 6, and 8. Suppose a sample size of 3 have to be drawn from it for the first, second and third rank;

1
Expert's answer
2021-11-01T19:52:06-0400

a.

 We have population values "1,3,4,6,8," population size "N=5"

"\mu=\dfrac{1+3+4+6+8}{5}=4.4"

b.


"\sigma^2=\dfrac{1}{5}((1-4.4)^2+(3-4.4)^2+(4-4.4)^2"

"+(6-4.4)^2)+(8-4.4)^2)=5.84"


c.


"\sigma=\sqrt{\sigma^2}=\sqrt{5.84}\approx2.4166"


d. We have population values "1,3,4,6,8" population size "N=5" and sample size "n=3." Thus, the number of possible samples which can be drawn without replacement is



"\dbinom{N}{n}=\dbinom{5}{3}=10""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 1,3,4 & 8/3 \\
 \hdashline
 2 & 1,3,6 & 10/3 \\
 \hdashline
 3 & 1,3,8 & 4 \\
 \hdashline
 4 & 1,4,6 & 11/3 \\
 \hdashline
 5 & 1,4,8 & 13/3\\
 \hdashline
 6 & 1,6,8 & 5 \\
 \hdashline
 7 & 3,4,6 & 13/3 \\
 \hdashline
 8 & 3,4,8 & 5 \\
 \hdashline
 9 & 3,6,8 & 17/3 \\
 \hdashline
 10 & 4,6,8 & 6 \\
 \hline
\end{array}"

The sampling distribution of the sample means.



"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c:c}
& \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
 & 8/3 & 1 & 1/10 & 8/30 & 64/90 \\
 \hdashline
 & 10/3 & 1 & 1/10 & 10/30 & 100/90 \\
 \hdashline
 & 11/3 & 1 & 1/10 & 11/10 & 121/90 \\
 \hdashline
 & 4 & 1& 1/10 & 12/30 & 144/90 \\
 \hdashline
 & 13/3 & 2 & 2/10 & 26/30 & 338/90 \\
 \hdashline
 & 5 & 2 & 2/10 & 30/30 & 450/90 \\
 \hdashline
 & 6 & 1 & 1/10 & 18/30 & 324/90 \\
 \hdashline
 & 17/3 & 1 & 1/10 & 17/30 & 289/90 \\
 \hdashline
 Total & & 10 & 1 & 4.4 & 61/3 \\ \hline
\end{array}"




"E(\bar{X})=\sum\bar{X}f(\bar{X})=4.4"


The mean of the sampling distribution of the sample means is equal to the the mean of the population.


"E(\bar{X})=\mu_{\bar{X}}=4.4=\mu"



e.


"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=\dfrac{61}{3}-(4.4)^2=\dfrac{2.92}{3}"




"\sigma_{\bar{X}}=\sqrt{Var(\bar{X})}=\sqrt{\dfrac{2.92}{3}}\approx0.9866"

Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{5.84}{3}(\dfrac{5-3}{5-1})"




"=\dfrac{2.92}{3}, True"